# Enhancing Job Salary Prediction with Disentangled Composition Effect Modeling: A Neural Prototyping Approach

### v1

@@ 5.2. Overall Performance Evaluation
The overall prediction performance of LGDESetNet and all baselines on four datasets is reported in Table 3. We can observe that: (1) LGDESetNet consistently outperforms existing salary prediction baselines across all datasets, demonstrating that our comprehensive framework more effectively captures the complex factors influencing salary. This validates our hypothesis that a holistic skill set-based reasoning process can effectively model intrinsic salary-influential patterns. (2) Compared to self-explainable baselines, LGDESetNet achieves an average improvement of 10%. This substantial gain stems from our model’s ability to disentangle and prototype relationships among set entities while maintaining job context awareness, enabling more accurate identification of salary-influential skill patterns. (3) The notable underperformance of Set-Tree compared to other methods highlights the importance of our approach. Set-Tree’s limitations in capturing context-dependent nuances in set data and modeling combinatorial interactions among set elements underscore the challenges in explainable skill set modeling. (4) Supplementary experiments in Table 6 show that LGDESetNet achieves slightly better performance than complex black-box models. This demonstrates that our model can bridge the typical performance gap between interpretable and black-box methods, providing a more transparent and accurate set modeling solution.
(a) IT
(b) Designer
(c) High-tech
(d) Finance
Figure 5. Ablation studies on IT, Designer, High-tech, and Finance datasets.
5.3. Ablation Study
@@
We introduce three variants of our LGDESetNet by disabling some parts of the network to evaluate the effectiveness of each component: (1) w/o prot, which disables the prototypical set learning module; (2) w/o sub, which disables the disentangled discrete subset selection module; and (3) w/o rel, which utilizes a conventional prototype replacement strategy (Ming et al., 2019) instead of our embedding-projection training algorithm. The results are reported in Figure 5. It can be observed that removing any modules from LGDESetNet reduces its performance, as the disentangled subset selection and prototypical learning modules are essential for capturing the full utility of set-valued reasoning. Particularly, the removal of the disentangled subset selection module leads to the most significant performance drop, underscoring the importance of skill semantic disentanglement. LGDESetNet outperforms these two components, which indicates that it not only raises the model explainability but also can increase the fitting ability by considering semantic interactions between subsets. Furthermore, LGDESetNet shows a 2%-5% performance gain over its variant with a classical projection strategy (Ming et al., 2019), validating our embedding-projection training’s effectiveness in advancing prototype learning.
(a) Prototype Number $M$
(b) View Number $H$
@@
Figure 6. Parameter sensitivity of IT job postings.
5.4. Parameters Sensitivity
@@
We further conduct the hyper-parameter analysis of the prototype number $M$ and the view number $H$. Specifically, we evaluate the performance of LGDESetNet with varying $M$ and $H$ on the IT dataset, compared with instance-wise selection model SESM (Geng et al., 2020) and prototype-based model TesNet (Wang et al., 2021). The results are shown in Figure 6. With the number of prototypes $M$ increases, the performance gets minorly better and stabilizes within a certain range. This pattern suggests that our model effectively captures the intrinsic dimensionality of the skill embedding space with a relatively small number of prototypes. Importantly, LGDESetNet maintains consistent performance advantages over baseline methods across all prototype settings, highlighting its inherent robustness to this hyperparameter. The stability plateau indicates that excessive prototypes may not necessarily improve the model’s performance, as they could introduce noise and redundancy. By default, we choose $M=64$ in order to get more diverse and representative skill prototypes.
@@
The view number $H$ (Figure 6(b)) exhibits more pronounced sensitivity, with a steep performance improvement from $H=1$ to $H=4$, followed by marginal gains thereafter. This pattern reveals that multiple views are essential for capturing the multi-faceted nature of skill representations, but only up to a certain threshold. The rapid early improvement suggests that different views effectively capture complementary aspects of skill semantics. However, the diminishing returns beyond $H=4$ indicate a saturation point where additional views contribute minimal new information. Moreover, excessive view fragmentation can partition the semantic space too finely, potentially compromising both computational efficiency and interpretability. Based on this analysis, we adopt $H=4$ for our experiments as the optimal balance between model performance and semantic coherence.
Table 4. Results on subset cohesion score.
@@
5.5. Quantitative Explanation Analysis
We evaluate LGDESetNet’s ability to identify semantically cohesive skill subsets using the Subset Cohesion Score (SCS). SCS quantifies the semantic tightness of a subset $S$, calculated as $\text{SCS}(S)=\frac{2}{|S|(|S|-1)}\sum_{x_{i},x_{j}\in S}\mathcal{F}(x_{i},x_{j})$, where $\mathcal{F}$ indicates skill co-occurrence frequency across the dataset. As shown in Table 4, LGDESetNet substantially outperforms SESM across all four domains. These results demonstrate that our approach more effectively captures domain-specific skill relationships, with the greatest advantages observed in highly specialized fields where skill coherence is crucial. The consistent performance improvement across diverse domains indicates that LGDESetNet’s co-occurrence graph density regularization and global prototype-based selection mechanism successfully identify meaningful skill combinations rather than merely selecting individually relevant skills.
@@ 5.5. Quantitative Explanation Analysis
Figure 7. t-SNE visualization of skill set embeddings with different view numbers $H$.
5.6. View Visualization
Figure 7 shows the t-SNE visualization of extracted skill subsets with different view numbers $H$. With $H$ = 2 or 4, we can observe that the skill embeddings are more compact and well-separated. Some overlap between views is reasonable, as certain skills share common traits, but the overall semantic meaning remains distinct. This indicates that LGDESetNet effectively captures the intrinsic skill relationships. When expanded to $H$ = 8, the skill embeddings become more dispersed, with less clear boundaries between different skill clusters. This suggests that excessive views may introduce redundancy, potentially compromising the model’s interpretability. This is also consistent with our previous experimental findings (shown in Figure 6(b)) that $H$ = 4 is the optimal balance.
@@ 5.6. View Visualization
Figure 8. Global salary-influential prototypical skill sets concerning different job contexts.
5.7. Salary-influential Skill Set Analysis
5.7.1. Representative Skill Sets
@@
5.7.2. Skill Set Influence Concerning Different Job Contexts
@@
Figure 9. Reasoning processes of LGDESetNet on the IT-related job posting. LGDESetNet identifies semantically meaningful and disentangled subsets from the input set and thereafter provides representative subsets for evidence comparison.
5.8. Case study
@@
5.9. User Study
Following the design rationale of (Rong et al., 2023), we conduct a user study to evaluate LGDESetNet’s explainability by accessing the user’s comprehension of the model’s reasoning process (Understanding), their confidence and comfort with the recommendations (Trust), and the intuitiveness and ease of use of provided explanations (Usability). Specifically, we recruited 53 postgraduate students and staffs with the background of computer science or engineering from different universities for evaluation. We compare our network to a full black box (HSBMF), where people are simply told “a predicted result produced by a trained machine learning model” in text form, and a representative prototyping model (ProtoPNet), where the learned prototypical skill sets are presented as explanations. For fair comparison, all the designed questions stem from each model’s inherent modeling procedure. As shown in Table 5, We can observe that self-explainable models increase user confidence by providing the evidence of model inference. It shows the necessity of our efforts to qualify the impacts of job skills. Moreover, traditional prototyping methods fail to provide semantically dense skill sets, which might confuse and confound users. In contrast, LGDESetNet can enhance user experience by offering clear, separate skill compositions and using prototypical examples as comparative evidence.
Table 5. User study results: 5-point Likert scale ratings of three measured features for explanatory methods.
Methods HSBMF ProtoPNet LGDESetNet Understanding 3.742 3.927 4.421 Trust 3.385 3.521 4.220 Usability 3.046 4.028 4.623
6. Conclusion
@@ Appendix A Proof of Permutation Invariance of LGDESetNet
which implies $P\text{Sigmoid}(X)P^{T}=\text{Sigmoid}(PXP^{T})$. ∎
@@
Proof.
Considering the GSS function only adds element-wise operations on Sigmoid, it does not change the original permutation property of Sigmoid. ∎
Proposition 1.
Proof.
For the dot-product attention mechanism (Vaswani et al., 2017), we have $V_{h}^{\text{Q}}=V_{X}W^{\text{Q}}_{h}$, $V_{h}^{\text{K}}=V_{X}W^{\text{K}}_{h}$, and $V_{h}^{\text{V}}=V_{X}W^{\text{V}}_{h}$. Therefore, we can get
which implies the multi-view subset selection network is permutation-invariant by reason that the pooling function satisfies the property of permutation invariance. ∎
Proposition 2.
$\operatorname{LGDESetNet}(PV_{X})=\operatorname{LGDESetNet}(V_{X})$
Proof.
Note that in LGDESetNet, we append a prototypical part network to the multi-view subset selection network. This network compares the extracted subsets with prototypes for weighted similarity aggregation. Since the prototyping operation is directly added to each subset embeddings $Z^{\text{s}}_{h}=\mathcal{T}\left(E^{\text{s}}_{h}\right)$, it does not change the permutation property of the multi-view subset selection network. Consequently, we can get that our LGDESetNet satisfies the permutation invariance condition. ∎
Appendix B Additional Experiments
B.1. Detailed Dataset Characteristics
@@
Figure 10. Dataset statistics and analysis of frequent skill sets (minSup $\geq$ 0.01).
Figure 11. Reasoning process of LGDESetNet on Designer job posting.
B.2. Additional Case Study
Table 6. Additional performance evaluation of state-of-the-art set-based black-box models. LGDESetNet achieves comparable performance to the best-performing black-box baselines, with slight improvements. The best results are highlighted in bold.
Methods Skill-Salary Job Postings (Sun et al., 2021) Job Salary Benchmarking (Meng et al., 2022a)

### v3

@@ 5.2 Overall Performance Evaluation
The overall prediction performance of LGDESetNet and all baselines on four datasets is reported in Table 3. We can observe that: (1) LGDESetNet consistently outperforms existing salary prediction baselines across all datasets, demonstrating that our comprehensive framework more effectively captures the complex factors influencing salary. This validates our hypothesis that a holistic skill set-based reasoning process can effectively model intrinsic salary-influential patterns. (2) Compared to self-explainable baselines, LGDESetNet achieves an average improvement of 10%. This substantial gain stems from our model’s ability to disentangle and prototype relationships among set entities while maintaining job context awareness, enabling more accurate identification of salary-influential skill patterns. Unlike existing self-interpretable models that often trade accuracy for explainability, LGDESetNet matches or even exceeds the performance of complex black-box models, as shown in Supplementary Table 6. This demonstrates its ability to bridge the typical performance gap between interpretable and black-box approaches, offering a more transparent yet accurate solution for set modeling. (3) The notable underperformance of Set-Tree compared to other methods highlights the importance of our approach. Set-Tree’s limitations in capturing job context-dependent nuances in set data and modeling combinatorial interactions among set elements underscore the challenges in explainable skill-based set modeling.
5.3 Ablation Study
@@
We introduce three variants of our LGDESetNet by disabling some parts of the network to evaluate the effectiveness of each component: (1) w/o prot, which disables the prototypical set learning module; (2) w/o sub, which disables the disentangled discrete subset selection module; and (3) w/o rel, which utilizes a conventional prototype replacement strategy [47] instead of our embedding-projection training algorithm. The results are reported in Figure 5. It can be observed that removing any modules from LGDESetNet reduces its performance, as the disentangled subset selection and prototypical learning modules are essential for capturing the full utility of set-valued reasoning. Particularly, the removal of the disentangled subset selection module leads to the most significant performance drop, underscoring the importance of skill semantic disentanglement. LGDESetNet outperforms these two components, which indicates that it not only raises the model explainability but also can increase the fitting ability by considering semantic interactions between subsets. Furthermore, LGDESetNet shows a 2%-5% performance gain over its variant with a classical projection strategy [47], validating our embedding-projection training’s effectiveness in advancing prototype learning.
@@
Figure 6: Parameter sensitivity of IT job postings. Left: Prototype number $M$. Right: View number $H$.
5.4 Parameters Sensitivity
@@
We further conduct the hyper-parameter analysis of the prototype number $M$ and the view number $H$. Specifically, we evaluate the performance of LGDESetNet with varying $M$ and $H$ on the IT dataset, compared with instance-wise selection model SESM [59] and prototype-based model TesNet [44]. The results are shown in Figure 6. With the number of prototypes $M$ increases, the performance gets minorly better and stabilizes within a certain range. This pattern suggests that our model effectively captures the intrinsic dimensionality of the skill embedding space with a relatively small number of prototypes. Importantly, LGDESetNet maintains consistent performance advantages over baseline methods across all prototype settings, highlighting its inherent robustness to this hyperparameter. The stability plateau indicates that excessive prototypes may not necessarily improve the model’s performance, as they could introduce noise and redundancy. By default, we choose $M=64$ in order to get more diverse and representative skill prototypes.
@@
The view number $H$ (Figure 6) exhibits more pronounced sensitivity, with a steep performance improvement from $H=1$ to $H=4$, followed by marginal gains thereafter. This pattern reveals that multiple views are essential for capturing the multi-faceted nature of skill representations, but only up to a certain threshold. The rapid early improvement suggests that different views effectively capture complementary aspects of skill semantics. However, the diminishing returns beyond $H=4$ indicate a saturation point where additional views contribute minimal new information. Moreover, excessive view fragmentation can partition the semantic space too finely, potentially compromising both computational efficiency and interpretability. Based on this analysis, we adopt $H=4$ for our experiments as the optimal balance between model performance and semantic coherence.
Table 4: Results on subset cohesion score.
@@
5.5 Quantitative Explanation Analysis
We evaluate LGDESetNet’s ability to identify semantically cohesive skill subsets using the Subset Cohesion Score (SCS). SCS quantifies the semantic tightness of a subset $S$, calculated as $\text{SCS}(S)=\frac{2}{|S|(|S|-1)}\sum_{x_{i},x_{j}\in S}\mathcal{F}(x_{i},x_{j})$, where $\mathcal{F}$ indicates skill co-occurrence frequency across the dataset. As shown in Table 4, LGDESetNet substantially outperforms SESM across all four domains. These results demonstrate that our approach more effectively captures domain-specific skill relationships, with the greatest advantages observed in highly specialized fields where skill coherence is crucial. The consistent performance improvement across diverse domains indicates that LGDESetNet’s co-occurrence graph density regularization and global prototype-based selection mechanism successfully identify meaningful skill compositions rather than merely selecting individual skills.
@@ 5.5 Quantitative Explanation Analysis
Figure 7: t-SNE visualizations of skill set embeddings under different numbers of views $H$.
5.6 View Visualization
Figure 7 shows the t-SNE visualization of extracted skill subsets with different view numbers $H$. With $H$ = 2 or 4, we can observe that the skill embeddings are more compact and well-separated. Some overlap between views is reasonable, as certain skills share common traits, but the overall semantic meaning remains distinct. This indicates that LGDESetNet effectively captures the intrinsic skill relationships. When expanded to $H$ = 8, the skill embeddings become more dispersed, with less clear boundaries between different skill clusters. This suggests that excessive views may introduce redundancy, potentially compromising the model’s interpretability. This is also consistent with our previous experimental findings (shown in Figure 6) that $H$ = 4 is the optimal balance.
@@ 5.6 View Visualization
Figure 8: Global salary-influential prototypical skill sets concerning different job contexts.
Figure 9: Reasoning processes of LGDESetNet on the IT-related job posting. LGDESetNet identifies semantically meaningful and disentangled subsets from the input set and thereafter provides representative subsets for evidence comparison.
5.7 Salary-influential Skill Set Analysis
5.7.1 Representative Skill Sets
@@
5.7.2 Skill Set Influence Concerning Different Job Contexts
@@
5.8 Case study
@@
Table 5: User study results: 5-point Likert scale ratings of three measured features for explanatory methods.
Methods Black box ProtoPNet LGDESetNet Understanding 3.742 3.927 4.421 Trust 3.385 3.521 4.220 Usability 3.046 4.028 4.623
5.9 User Study
Following the design rationale of [64], we conduct a user study to evaluate LGDESetNet’s explainability by accessing the user’s comprehension of the model’s reasoning process (Understanding), their confidence and comfort with the recommendations (Trust), and the intuitiveness and ease of use of provided explanations (Usability). Specifically, we recruited 53 postgraduate students and staffs with the background of computer science or engineering from different universities for evaluation. We compare our network to a full black box (HSBMF), where people are simply told “a predicted result produced by a trained machine learning model” in text form, and a representative prototyping model (ProtoPNet), where the learned prototypical skill sets are presented as explanations. For fair comparison, all the designed questions stem from each model’s inherent modeling procedure. As shown in Table 5, We can observe that self-explainable models increase user confidence by providing the evidence of model inference. It shows the necessity of our efforts to qualify the impacts of job skills. Moreover, traditional prototyping methods fail to provide semantically dense skill sets, which might confuse and confound users. In contrast, LGDESetNet can enhance user experience by offering clear, separate skill compositions and using prototypical examples as comparative evidence.
6 Conclusion
@@
{acknowledgement}
This work is partly supported by the National Key Research and Development Program of China (No. 2023YFF0725001), the National Natural Science Foundation of China (No. 62306255, 92370204, 62176014), the Natural Science Foundation of Guangdong Province (No. 2024A1515011839), the Fundamental Research Project of Guangzhou (No. 2024A04J4233), the Guangzhou-HKUST(GZ) Joint Funding Program (No.2023A03J0008), and the Education Bureau of Guangzhou Municipality.
Appendixes
@@ Appendix A Proof of Permutation Invariance of LGDESetNet
which implies $P\text{Sigmoid}(X)P^{T}=\text{Sigmoid}(PXP^{T})$.
@@
Proof Appendix A.5.
Considering the GSS function only adds element-wise operations on Sigmoid, it does not change the original permutation property of Sigmoid.
@@
Figure 10: Dataset statistics and analysis of frequent skill sets (minSup $\geq$ 0.01).
Figure 11: Reasoning process of LGDESetNet on Designer job posting.
Table 6: Additional performance evaluation of SOTA set-based black-box models. LGDESetNet achieves comparable performance to the best-performing black-box baselines, with slight improvements. The best results are highlighted in bold.
Methods Skill-Salary Job Postings [4] Job Salary Benchmarking [8]
@@
Proposition 1.
Proof Appendix A.6.
For the dot-product attention mechanism [57], we have $V_{h}^{\text{Q}}=V_{X}W^{\text{Q}}_{h}$, $V_{h}^{\text{K}}=V_{X}W^{\text{K}}_{h}$, and $V_{h}^{\text{V}}=V_{X}W^{\text{V}}_{h}$. Therefore, we can get
which implies the multi-view subset selection network is permutation-invariant by reason that the pooling function satisfies the property of permutation invariance.
Proposition 2.
$\operatorname{LGDESetNet}(PV_{X})=\operatorname{LGDESetNet}(V_{X})$
Proof Appendix A.7.
Note that in LGDESetNet, we append a prototypical part network to the multi-view subset selection network. This network compares the extracted subsets with prototypes for weighted similarity aggregation. Since the prototyping operation is directly added to each subset embeddings $Z^{\text{s}}_{h}=\mathcal{T}\left(E^{\text{s}}_{h}\right)$, it does not change the permutation property of the multi-view subset selection network. Consequently, we can get that our LGDESetNet satisfies the permutation invariance condition.
Appendix B Additional Experiments
Appendix B.1 Detailed Dataset Characteristics
Appendix B.2 Additional Case Study